40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Oct 26 23:03:03 2011, kiwamu, Update, LSC, RF distribution box : two more 55MHz available 
    Reply  Thu Oct 27 16:09:37 2011, kiwamu, Update, LSC, RF distribution box : two more 55MHz available RF_Distribution_Box.png
Message ID: 5744     Entry time: Wed Oct 26 23:03:03 2011     Reply to this: 5746
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: RF distribution box : two more 55MHz available 

The RF distribution box has been modified so that it generates two more 55 MHz LO signals.

After the modification I put the box back in place.

Then I checked the MICH and YARM locking quickly as a working test of the distribution box and it is working fine so far.

I will update the diagram of the RF distribution box (#4342) tomorrow.

 

(Motivation)

  Since we newly installed POP55 (#5743) an LO signal was needed for the demodulation.

However the RF distribution box didn't have any extra LO outputs.

Therefore we had to make a modification on the RF distribution box so that we can have a 55MHz LO signal for POP55.

Eventually I made two more 55MHz outputs including one spare.

 

(Modification)

The box actually had two extra output SMAs which had been just feed-thru connectors on the front panel without any signals going through.

In the box the modules consist of two categories; the 11MHz system and 55MHz system. I modified only a part of the 55MHz system.

The modification was done in this way:

  * split two branches of 55MHz into four branches by installing two new power splitters (ZMSC-2-1).

  * made and installed some SMA cables whose length were adjusted to be nicely fit in the box.

  * readjustment of the RF levels to 0-2 dBm at the outputs by replacing some attenuators.

  * checked the signals if all of them were happily coming out or not.

Also I found that the POX11 and POY11 demod boards were connected to the whitening filters in a wrong way.

The I and Q signals were in a wrong order. So I corrected them so that the upper inputs in the whitening filter is always the I signal.

 

(RF levels on 55MHz LO outputs)

Since the demod board requires a certain level of the RF signal as as LO, the LO signals have to be 0-2 dBm.

Here are the RF level in each 55MHz output after the adjustment of the level.

     AS55 =  0.76 dBm

     REFL55 = 0.76 dBm

     POP55 = 0.79 dBm

     spare = 0.83 dBm

Those numbers were measured by an oscilloscope and the oscilloscope was configured to measure the rms with the input impedance matched to 50Ohm.

In the measurement I used the actual input seed 55MHz signal from the RF generation box to drive the distribution box.

ELOG V3.1.3-